A board, 44 cm long is cut into three pieces such that the second piece is twice as long as the first and the third is 4 cm longer than the second.
Find the length of the shorter piece.

Answer by algebrahouse.com(1659)

x = length of first piece
2x = length of second piece {second piece is twice as long as the first}
2x + 4 = length of third piece {third piece is 4 cm longer than the second}

x + 2x + 2x + 4 = 44 {total length of all pieces is 44}
5x + 4 = 44 {combined like terms}
5x = 40 {subtracted 4 from each side}
x = 8 {divided each side by 5}

shorter piece is 8 cm

For more help from me, visit: www.algebrahouse.com